# Einstein from Noise: Statistical Analysis

### v2

@@ A.3 Conditioning on the Fourier frequency noise component
Let $\mathsf{S}^{(+)}\sim{\cal N}\left(\mu,\mathsf{\Sigma}\right)$ and $\mathsf{S}^{(-)}\sim{\cal N}\left(-\mu,\mathsf{\Sigma}\right)$ be two $d$-dimensional Gaussian vectors, where $\mathsf{\Sigma}$ is a circulant covariance matrix, with $\mathsf{rank}(\mathsf{\Sigma})>d/2$. We define the entries of $\mu$ as,
for $\varphi\in[0,2\pi)$, and $10\leq\ell\leq d-$ 1. Note that $-\mu_{\ell}=\cos{\left(\frac{2\pi k}{d}\ell+\varphi+\pi\right)}$, for $10\leq\ell\leq d-$ 1. Define,
@@
Consider the definitions in (A.34)–(A.36), and assume $\mathsf{rank}(\mathsf{\Sigma})>d/2$. Fix $0\leq\ell\leq{d-1}$. If $\mu_{\ell}>0$, then,
@@
$\mathbb{P}\left[\hat{\mathsf{R}}^{(-)}=\ell\right]<\mathbb{P}\left[\hat{\mathsf{R}}^{(+)}=\ell\right].$ (A.38)
In particular, for any $\varphi\in[0,2\pi)$ and $10\leq k\leq d-$ 1,
@@
for $10\leq\ell\leq d-$ 1. Since $\mathsf{S}^{(+)}$ and $\mathsf{S}^{(-)}$ can be decomposed as $\mathsf{S}^{(+)}=\mathsf{Z}+\mu$ and $\mathsf{S}^{(-)}=\mathsf{Z}-\mu$, where $\mathsf{Z}$ is a cyclo-stationary process, and $\mu$ is defined in (A.34). Then,
@@
By definition, since $\mathsf{Z}$ is a zero-mean Gaussian, cyclo-stationary random process, its cumulative distribution function $F_{\mathsf{Z}}$ is invariant under cyclic shifts, i.e.,
@@
for $11\leq m\leq d-$ 1. It is clear from (A.47) that $\mathsf{Z}^{(1)}$ and $\mathsf{Z}^{(2)}$ have the same cumulative distribution function, i.e.,
@@
where the second equality follows from (A.50). Next, we note that for every $10<m\leq d-$ 1 and $\mu_{\ell}>0$,
@@
According to (A.53), the inequalities in (A.54) and (A.55) are strict whenever $\cos\left(\frac{2\pi k}{d}m\right)<1$, which holds for the majority of values of $m$. In particular, at least $d/2$ of the inequalities are strict for $10\leq m\leq d-$ 1.
@@
Assume $\mathsf{\mathrm{rank}}\left(\Sigma\right)>d/2$. Then, for $\mu_{\ell}>0$, we have,
@@
By (A.37)–(A.38), as for any $10\leq\ell\leq d-$ 1 such that $\mu_{\ell}=\cos\left(\frac{2\pi k}{d}\ell+\varphi\right)>0$ it holds that $\mathbb{P}[\hat{\mathsf{R}}^{(+)}=\ell]>\mathbb{P}[\hat{\mathsf{R}}^{(-)}=\ell]$, otherwise, for $10\leq\ell\leq d-$ 1 such that $\mu_{\ell}=\cos\left(\frac{2\pi k}{d}\ell+\varphi\right)<0$, it holds that $\mathbb{P}[\hat{\mathsf{R}}^{(+)}=\ell]<\mathbb{P}[\hat{\mathsf{R}}^{(-)}=\ell]$. Therefore,
@@
=$\mathcal{M}=\left\{m:\mu_{\ell}-\mu_{\ell+m}>\mu_{\ell-m}+\mu_{\ell}\right\}.$ (A.65)
@@
i.e., the event that $\mathsf{Z}_{\ell-r}-\mu_{\ell-r}+\mu_{\ell}$ attains the maximum in the expression above. Since $\mathsf{rank}(\Sigma)>d/2$ and the number of strict inequalities in (A.65) is $\left|\mathcal{M}\right|\geq d/2$, by the pigeonhole principle, there exists $r\neq\ell$ such that the next both conditions are satisfied,
$r\in\mathcal{M},\quad and,\quad\mathbb{P}\left[{\mathcal{C}_{r}}\right]=\mathbb{P}\left[\max_{m\neq 0}\left\{\mathsf{Z}_{\ell-m}-\mu_{\ell-m}+\mu_{\ell}\right\}=\left\{\mathsf{Z}_{\ell-r}-\mu_{\ell-r}+\mu_{\ell}\right\}\right]>0.$ (A.67)
@@
which proves (A.56), and completes the proof.
Appendix B Auxiliary statements for Theorem 4.1
In this section, we prove several auxiliary statements needed in the proof of Theorem 4.1. Recall the definition of $\phi_{e,i}[k]$ in (A.7) and of $\mu_{\mathsf{A},k},\mu_{\mathsf{B},k},\sigma_{\mathsf{A},k}^{2},\sigma_{\mathsf{B},k}^{2}$ in (A.14)–(A.17).
@@
$\mathsf{\hat{X}}[k]$ ${\mathsf{a.s.}}e^{j\phi_{\mathsf{X}}[k]}\cdot\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right]$ (B.4)
@@ B.2 Convergence to non-vanishing signal
Proposition B.2.
Recall the definition of $\phi_{e,i}[k]$ in (A.7). Fix $d\in\mathbb{N}$, and assume that $\mathsf{X}[k]\neq 0$ for all $10<k\leq d-$ 1. Then, for any $10\leq k\leq d-$ 1,
@@
where the right-hand side is defined as in (A.18). By (A.19), the mean vector of $\mathsf{S}_{1}|\mathsf{N}_{1}[k]$ has a cosine trend, as assumed in Proposition A.3 in (A.34). Additionally, $\mathsf{S}_{1}|\mathsf{N}_{1}[k]$ is a Gaussian cyclo-stationary process, as described in (A.20). The final condition to verify is that the covariance matrix of $\mathsf{S}_{1}|\mathsf{N}_{1}[k]$ satisfies $\mathsf{rank}(\mathsf{\Sigma})>d/2$.
By Proposition B.2, we assume $\mathsf{X}[k]\neq 0$ for all $10<k\leq d-$ 1, which implies $\mathsf{rank}(\mathsf{\Sigma})>d/2$. Indeed, for the covariance matrix to have rank greater than $d/2$, at least half of its eigenvalues must be nonzero. As proved in Lemma A.1, the eigenvalues of $\mathsf{\Sigma}$ are given by $|\tilde{\mathsf{X}}_{k}[\ell]|^{2}$ for $10\leq\ell\leq d-$ 1. Since the spectrum of $\mathsf{X}$ is non-vanishing, it follows that $\mathsf{rank}(\mathsf{\Sigma})>d/2$. In fact, we obtain $\mathsf{rank}(\mathsf{\Sigma})\geq d-2$, which is larger than $d/2$ for $d>4$.
Finally, since the conditional distribution of $\hat{\mathsf{R}}_{1}$ given $\{\left|\mathsf{N}_{1}[k]\right|=n,\phi_{\mathsf{N}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi\}$ matches that of $\hat{\mathsf{R}}^{(+)}$ in (A.35), and similarly, given $\{\left|\mathsf{N}_{1}[k]\right|=n,\phi_{\mathsf{N}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi+\pi\}$, it matches $\hat{\mathsf{R}}^{(-)}$ in (A.36), the sum of the integrands on the right-hand side of (B.15) equals the left-hand side of (A.39). By Proposition A.3, this sum is positive for all $\varphi\in[0,\pi]$. Together with (B.14), this completes the proof of Proposition B.2.
@@ B.3 Convergence rate in distribution of the Fourier phases
Proposition B.3.
Fix $d\in\mathbb{N}$. Then, for any $10\leq k\leq d-$ 1,
@@ B.4 Convergence rate in expectation of the Fourier phases
Proposition B.4.
Recall the definitions of $\mu_{\mathsf{B},k}$, and $\sigma_{\mathsf{A},k}^{2}$ in (A.15), and (A.17), respectively. Assume that $\mathsf{X}[k]\neq 0$, for all $10<k\leq d-$ 1. Then, as $M\to\infty$,
@@
for every $10\leq k\leq d-$ 1. Recall by the definition of $\mathsf{Q}_{k}$ in (B.3) that $\mathbb{E}\left[\mathsf{Q}_{k}^{2}\right]=\sigma_{\mathsf{A},k}^{2}/\mu_{\mathsf{B},k}^{2}$, which is equivalent to the right-hand-side of (B.24).
@@
Let us denote the summand in the denominator in (B.26) by $V_{i}\triangleq\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)$, for $10\leq i\leq M-$ 1. Then, we note that,
@@
Thus, the second term at the r.h.s. of (B.27) indeed converges to zero as $1/M^{2}$.
B.4.2 Proof of Lemma B.6
@@
Since $[\arctan(x)]^{2}/x^{2}\to 1$ as $x\to 0$, it follows that the Taylor expansion of $\arctan(x)$ around $x=0$, which holds for $\left|x\right|<1$, and is applicable on the event $\left\{\mathsf{A}_{M}<\epsilon\sqrt{M}\right\}$:
@@
Since (B.56) is true for every $\epsilon_{2}>0$, we get the lower bound in (B.40), which concludes the proof of (B.41).
Appendix C Proof of Theorem 4.1
@@
$\left|\mathsf{\hat{X}}[k]e^{-j\phi_{\mathsf{X}}[k]}\right|\xrightarrow[]{\mathsf{a.s.}}\Bigl{\lvert}\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right]\Bigr{\rvert},$ (C.1)
@@
$\mu_{\mathsf{A},k}=\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right]=0.$ (C.2)
@@
$\mu_{\mathsf{B},k}=\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]>0.$ (C.3)
Combining (C.1), (C.2), and (C.3) proves the convergence of the estimator’s magnitudes of (4.3).
@@
$\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=\arctan\left(\frac{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\right),$ (C.4)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}0.$ (C.5)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}\frac{\mu_{\mathsf{A},k}}{\mu_{\mathsf{B},k}},$ (C.6)
@@
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\frac{\sigma_{\mathsf{A},k}^{2}}{\mu_{\mathsf{B},k}^{2}},$ (C.7)
@@
$C_{k}\triangleq\frac{\sigma_{\mathsf{A},k}^{2}}{\mu_{\mathsf{B},k}^{2}}=\frac{\mathbb{E}\left(\left[\left|\mathsf{N}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]^{2}\right)}{\left(\mathbb{E}{\left[\left|\mathsf{N}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]}\right)^{2}}.$ (C.8)
Appendix D High-dimensional signal convergence

### v3

@@ A.5 Positive probability of each maximizer event
Let $\mathsf{S}^{(+)}\sim{\cal N}\!\left(\mu,\mathsf{\Sigma}\right)$ and $\mathsf{S}^{(-)}\sim{\cal N}\!\left(-\mu,\mathsf{\Sigma}\right)$ be two $d$-dimensional Gaussian vectors, where $\mathsf{\Sigma}$ is a real, symmetric, circulant covariance matrix. Define the maximizers
@@
and assume they are unique almost surely. We define the entries of $\mu$ as,
for $\varphi\in[0,2\pi)$, and $0\leq\ell\leq d-1$. Note that $\mu_{\ell}=\cos{\left(\frac{2\pi k}{d}\ell+\varphi+\pi\right)}$, for $0\leq\ell\leq d-1$.
@@
Consider the definitions in (A.59)–(A.58), and assume the maximizers in (A.58) are unique a.s. Moreover, assume that for every $r\in\{1,\ldots,d-1\}$,
$\mathbb{P}\!\left(\mathsf{S}^{(+)}_{r}>\max_{t\neq r}\mathsf{S}^{(+)}_{t}\right)>0\qquad\text{and}\qquad\mathbb{P}\!\left(\mathsf{S}^{(-)}_{r}>\max_{t\neq r}\mathsf{S}^{(-)}_{t}\right)>0.$ (A.60)
Fix $0\leq\ell\leq{d-1}$. If $\mu_{\ell}>0$, then,
@@
$\mathbb{P}\left[\hat{\mathsf{R}}^{(-)}=\ell\right]>\mathbb{P}\left[\hat{\mathsf{R}}^{(+)}=\ell\right].$ (A.62)
In particular, for any $\varphi\in[0,2\pi)$ and $0\leq k\leq d-1$,
@@
for $0\leq\ell\leq d-1$. Since $\mathsf{S}^{(+)}$ and $\mathsf{S}^{(-)}$ can be decomposed as $\mathsf{S}^{(+)}=\mathsf{Z}+\mu$ and $\mathsf{S}^{(-)}=\mathsf{Z}-\mu$, where $\mathsf{Z}$ is a cyclo-stationary process, and $\mu$ is defined in (A.59). Then,
@@
By definition, since $\mathsf{Z}$ is a zero-mean Gaussian, cyclo-stationary random process (i.e., with a real, symmetric, circulant covariance matrix), its cumulative distribution function $F_{\mathsf{Z}}$ is invariant under cyclic shifts, i.e.,
@@
for $1\leq m\leq d-1$. It is clear from (A.71) that $\mathsf{Z}^{(1)}$ and $\mathsf{Z}^{(2)}$ have the same cumulative distribution function, i.e.,
@@
where the second equality follows from (A.74). Next, we note that for every $0<m\leq d-1$ and $\mu_{\ell}>0$,
@@
According to (A.77), the inequalities in (A.78) and (A.79) are strict whenever $\cos\left(\frac{2\pi k}{d}m\right)<1$, which holds for the majority of values of $m$. In particular, at least $d/2$ of the inequalities are strict for $0\leq m\leq d-1$.
@@
Assume the maximizers in (A.58) are unique a.s. and satisfy (A.60). Then, for $\mu_{\ell}>0$, we have,
@@
By (A.61)–(A.62), as for any $0\leq\ell\leq d-1$ such that $\mu_{\ell}=\cos\left(\frac{2\pi k}{d}\ell+\varphi\right)>0$ it holds that $\mathbb{P}[\hat{\mathsf{R}}^{(+)}=\ell]>\mathbb{P}[\hat{\mathsf{R}}^{(-)}=\ell]$, otherwise, for $0\leq\ell\leq d-1$ such that $\mu_{\ell}=\cos\left(\frac{2\pi k}{d}\ell+\varphi\right)<0$, it holds that $\mathbb{P}[\hat{\mathsf{R}}^{(+)}=\ell]<\mathbb{P}[\hat{\mathsf{R}}^{(-)}=\ell]$. Therefore,
@@
$\mathcal{M}=\left\{m:\mu_{\ell}-\mu_{\ell+m}>\mu_{\ell-m}-\mu_{\ell}\right\}.$ (A.89)
@@
i.e., the event that $\mathsf{Z}_{\ell-r}-\mu_{\ell-r}+\mu_{\ell}$ attains the maximum in the expression above. By assumption (A.60), the event $\mathcal{C}_{r}$ for every $r$, i.e. $\mathbb{P}(\mathcal{C}_{r})>0$. Thus, we may choose $r\in\mathcal{M}$ with $\mathbb{P}(\mathcal{C}_{r})>0$, i.e.,
$r\in\mathcal{M},\quad\text{and}\quad\mathbb{P}(\mathcal{C}_{r})=\mathbb{P}\!\left(\max_{m\neq 0}\Big\{\mathsf{Z}_{\ell-m}-\mu_{\ell-m}+\mu_{\ell}\Big\}=\mathsf{Z}_{\ell-r}-\mu_{\ell-r}+\mu_{\ell}\right)>0.$ (A.91)
@@
which proves (A.80), and completes the proof. ∎
Appendix B Proof of Theorem 4.1
First, we prove several auxiliary statements needed in the proof of Theorem 4.1. Recall the definition of $\phi_{e,i}[k]$ in (A.7) and of $\mu_{\mathsf{A},k},\mu_{\mathsf{B},k},\sigma_{\mathsf{A},k}^{2},\sigma_{\mathsf{B},k}^{2}$ in (A.14)–(A.17).
@@
$\mathsf{\hat{X}}[k]$ $\xrightarrow[]{\mathsf{a.s.}}e^{j\phi_{\mathsf{X}}[k]}\cdot\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right]$ (B.4)
@@ B.2 Convergence to non-vanishing signal
Proposition B.2 (Convergence to non-vanishing signal).
Recall the definition of $\phi_{e,i}[k]$ in (A.7). Fix $d\in\mathbb{N}$, and assume that $\mathsf{X}[k]\neq 0$ for all $0<k\leq d-1$. Then, for any $0\leq k\leq d-1$,
@@
where the right-hand side is defined as in (A.18). By (A.19), the mean vector of $\mathsf{S}_{1}|\mathsf{N}_{1}[k]$ has a cosine trend, as assumed in Proposition A.5 in (A.59). Additionally, $\mathsf{S}_{1}|\mathsf{N}_{1}[k]$ is a Gaussian cyclo-stationary process, as described in (A.20). The final condition to verify is (A.60), which is satisfied by Lemma A.4.
Since the conditional distribution of $\hat{\mathsf{R}}_{1}$ given $\{\left|\mathsf{N}_{1}[k]\right|=n,\phi_{\mathsf{N}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi\}$ matches that of $\hat{\mathsf{R}}^{(+)}$ in (A.57), and similarly, given $\{\left|\mathsf{N}_{1}[k]\right|=n,\phi_{\mathsf{N}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi+\pi\}$, it matches $\hat{\mathsf{R}}^{(-)}$ in (A.58), the sum of the integrands on the right-hand side of (B.15) equals the left-hand side of (A.63). By Proposition A.5, this sum is positive for all $\varphi\in[0,\pi]$. Together with (B.14), this completes the proof of Proposition B.2.
@@ B.3 Convergence rate in distribution of the Fourier phases
Proposition B.3 (Convergence in distribution of the Fourier phases).
Fix $d\in\mathbb{N}$. Then, for any $0\leq k\leq d-1$,
@@ B.4 Convergence rate in expectation of the Fourier phases
Proposition B.4 (Convergence rate of the Fourier phases).
Recall the definitions of $\mu_{\mathsf{B},k}$, and $\sigma_{\mathsf{A},k}^{2}$ in (A.15), and (A.17), respectively. Assume that $\mathsf{X}[k]\neq 0$, for all $0<k\leq d-1$. Then, as $M\to\infty$,
@@
for every $0\leq k\leq d-1$. Recall by the definition of $\mathsf{Q}_{k}$ in (B.3) that $\mathbb{E}\left[\mathsf{Q}_{k}^{2}\right]=\sigma_{\mathsf{A},k}^{2}/\mu_{\mathsf{B},k}^{2}$, which is equivalent to the right-hand-side of (B.24).
@@
Let us denote the summand in the denominator in (B.26) by $V_{i}\triangleq\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)$, for $0\leq i\leq M-1$. Then, we note that,
@@
Thus, the second term at the r.h.s. of (B.27) indeed converges to zero as $1/M^{2}$. ∎
Proof of Lemma B.6.
@@
Since $\arctan(x)]^{2}/x^{2}\to 1$ as $x\to 0$, it follows that the Taylor expansion of $\arctan(x)$ around $x=0$, which holds for $\left|x\right|<1$, and is applicable on the event $\left\{\mathsf{A}_{M}<\epsilon\sqrt{M}\right\}$:
@@
Since (B.56) is true for every $\epsilon_{2}>0$, we get the lower bound in (B.40), which concludes the proof of (B.41). ∎
B.5 Proof of Theorem 4.1
@@
$\left|\mathsf{\hat{X}}[k]e^{-j\phi_{\mathsf{X}}[k]}\right|\xrightarrow[]{\mathsf{a.s.}}\Bigl\lvert\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right]\Bigr\rvert$, (B.57)
@@
$\mu_{\mathsf{A},k}=\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right]=0.$ (B.58)
@@
$\mu_{\mathsf{B},k}=\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]>0.$ (B.59)
Combining (B.57), (B.58), and (B.59) proves the convergence of the estimator’s magnitudes of (4.3).
@@
$\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=\arctan\left(\frac{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\right),$ (B.60)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}0.$ (B.61)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{N}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}\frac{\mu_{\mathsf{A},k}}{\mu_{\mathsf{B},k}},$ (B.62)
@@
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\frac{\sigma_{\mathsf{A},k}^{2}}{\mu_{\mathsf{B},k}^{2}},$ (B.63)
@@
$C_{k}\triangleq\frac{\sigma_{\mathsf{A},k}^{2}}{\mu_{\mathsf{B},k}^{2}}=\frac{\mathbb{E}\left(\left[\left|\mathsf{N}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]^{2}\right)}{\left(\mathbb{E}{\left[\left|\mathsf{N}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]}\right)^{2}}.$ (B.64)
Appendix C High-dimensional argmax asymptotics